OpenStudy (anonymous):

What is the value of the y-coordinate of the solution to the system of equations? x+3y = 2 -3x-8y=4 Plese show steps on how to solve thanks!

OpenStudy (anonymous):

(X+3Y=2)*3 3X+6Y=6 -3X-8Y=4 add these equations your gona get -2Y=10 Y=-5

OpenStudy (anonymous):

How did you get that? what steps did you use? I've never done these before xD

OpenStudy (anonymous):

like why would you times everything by 3, i dont know.

OpenStudy (anonymous):

holda man

OpenStudy (anonymous):

just mulitply whole equation by 3 i mean the first one and then add both the equations

OpenStudy (anonymous):

to elimintate x

OpenStudy (anonymous):

but why 3?

OpenStudy (anonymous):

from my perspective, i dont know why we are doing that with 3? is it because 3 is the y cordinate? i dont know. so i am lost

OpenStudy (anonymous):

because in the second equation its -3x so we are equating by multiplying 3 in the first one and obtaining equal and opposite and hence eliminating x from both the parts

OpenStudy (anonymous):

still any doubts

OpenStudy (anonymous):

yep, when looking at the problem i am asking myself, what is the first thing i should do and why.

OpenStudy (anonymous):

see firstly u have to eliminate x in order to get the value of y right?

OpenStudy (anonymous):

for x elimination we are multiplying the first equation by 3

OpenStudy (anonymous):

3x+9y=6 -3x-8y=4 add the above two equations y=2

OpenStudy (anonymous):

sorry after addition y=10

OpenStudy (anonymous):

See nits, i understand yours because his said 3x+6y=6y

OpenStudy (anonymous):

so basically your just multiplying out the x, by 3 i knew there was something wrong with his example.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats why i was so confused, lol because 3 x 3 = 9 , where did he get 6 from . LOL

OpenStudy (anonymous):

yes.........any other doubts

OpenStudy (anonymous):

not doubts but question as to why we use that number 3, like if there was another problem. where did the 3 come from that we used?

OpenStudy (anonymous):

was the 3 from the lower problem with 3x?

OpenStudy (anonymous):

suppose.... 5x+4y=10 x+6y=9 then for eliminaton of y we would have multiplied 2nd equation by 5 5x+4y=10 5x+30y=45 and then subtract the above two equations which would give -26y=-35 y=35/26

OpenStudy (anonymous):

THank you , That is very clear and i understand it now =D

OpenStudy (anonymous):

Left handed brain, its stubborn to take in information that doesn't add up.